# DORSal: Diffusion for Object-centric

### v1

@@ Abstract
Recent progress in 3D scene understanding enables scalable learning of representations across large datasets of diverse scenes. As a consequence, generalization to unseen scenes and objects, rendering novel views from just a single or a handful of input images, and controllable scene generation that supports editing, is now possible. However, training jointly on a large number of scenes typically compromises rendering quality when compared to single-scene optimized models such as NeRFs. In this paper, we leverage recent progress in diffusion models to equip 3D scene representation learning models with the ability to render high-fidelity novel views, while retaining benefits such as object-level scene editing to a large degree. In particular, we propose DORSal, which adapts a video diffusion architecture for 3D scene generation conditioned on object-centric slot-based representations of scenes. On both complex synthetic multi-object scenes and on the real-world large-scale Street View dataset, we show that DORSal enables scalable neural rendering of 3D scenes with object-level editing and improves upon existing approaches.
${}^{\dagger}$ork done while interning at Google, $*{}^{*}$equal contribution.
@@ 1 Introduction
In this work, we combine techniques from both of these subfields to further neural 3D scene rendering. We leverage object-centric scene representations to condition probabilistic diffusion decoders capable of synthesizing novel views while also handling uncertainty about the scene. In particular, we use Object Scene Representation Transformer (OSRT) (Sajjadi et al., 2022a) to compute a set of Object Slots for a visual scene from only few observations, and condition a video diffusion architecture (Ho et al., 2022c) with these slots to generate sets of 3D consistent novel views of the same scene. We show that conditioning on object-level representations allows for scaling more gracefully to complex scenes, large sets of target views, and allows basic object-level scene editing by removing slots. In summary, our contributions are as follows:
  • We introduce Diffusion for Object-centric Representations of Scenes et al. (DORSal), an approach to 3D novel-view synthesis combining object representations with diffusion decoders.
  • Compared to prior methods from the 3D scene understanding literature (Sajjadi et al., 2022a, c), DORSal renders novel views that are significantly more precise (e.g. 5x-10x improvement in FID) while staying true to the content of the scene. Compared to prior work on 3D Diffusion Models (Watson et al., 2023), DORSal scales to more complex scenes, performing significantly better on real-world Street View data.
@@ 2 Preliminaries
Recent breakthroughs in neural rendering have inspired multiple works for 3D-centric object representations, including uORF (Yu et al., 2022) and ObSuRF (Stelzner et al., 2021). However, these methods do not scale beyond simple datasets due to the high memory and compute requirements of volumetric rendering. More recently, the Object Scene Representation Transformer (OSRT) (Sajjadi et al., 2022a) has been proposed as a powerful method that scales to much more complex datasets with wider camera pose distributions such as MultiShapeNet (Sajjadi et al., 2022c). Building upon SRT (Sajjadi et al., 2022c), it uses light-field rendering and to obtain speed-ups by a factor of $\mathcal{O}(100)$ at inference time.
Figure 1: Model overview. (a) OSRT is trained to predict novel views through an Encoder-Decoder architecture with an Object Slot latent representation of the scene. Since the model is trained with the L2 loss and the task contains significant amounts of ambiguity, the predictions are commonly blurry. (b) After training the OSRT model, we take the Object Slots and combine it with the target Poses to be used as conditioning. Our Multiview U-Net is trained in a diffusion process to denoise novel views while cross-attending into the conditioning features (see Figure 2 for details). This results in sharp renders at test time, which can still be decomposed into the objects in the scene to support edits.
@@
Once the input views are encoded into the Object Slots, arbitrary novel views can be rendered by passing the target ray origin and direction (the Pose) into the Decoder. To encourage an object-centric decomposition in the Object Slots, Spatial Broadcast Decoders Watters et al. (2019) are commonly used in the literature: Each slot is decoded independently into a pair of RGB and alpha using the same decoder, after which a Softmax over the slots decides on the final output color. Since OSRT is trained end-to-end with the L2 loss, any uncertainty about novel views necessarily leads to blur in the final renders.
@@
Denoising Diffusion Probabilistic Models (DDPMs) learn to generate data $\boldsymbol{x}$ by learning the reverse of a simple destruction process (Sohl-Dickstein et al., 2015). Such a diffusion process is convenient to express in its marginal form:
$q(\boldsymbol{z}_{t}|\boldsymbol{x})=\mathcal{N}(\boldsymbol{z}_{t}|\alpha_{t}\boldsymbol{x},\sigma_{t}^{2}\mathbf{I}),$ (1)
where $\alpha_{t}$ is a decreasing function and $\sigma_{t}$ is an increasing function over diffusion time $t\in[0,1]$. A neural network is then used to approximate $\boldsymbol{\epsilon}_{t}$, the reparametrization noise to sample $\boldsymbol{z}_{t}$:
$L=\mathbb{E}_{t\sim\mathcal{U}(0,1),\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,\mathbf{I})}\Big{[}w(t)||\boldsymbol{\epsilon}_{t}-f(\boldsymbol{z}_{t},t)||^{2}\Big{]},$ (2)
where $f$ is a neural network and $\boldsymbol{z}_{t}=\alpha_{t}\boldsymbol{x}+\sigma_{t}\boldsymbol{\epsilon}_{t}$. There exists a particular weighting $w(t)$ for this objective to be a variational negative lowerbound on $\log p(\boldsymbol{x})$, although in practice the constant weighting $w(t)=1$ has been found to be superior for sample quality Ho et al. (2020); Kingma et al. (2023). Because diffusion models learn to correlate the pixels in their generations, they are able to generate images with crisp details even if the exact location of such details is not entirely known.
A diffusion model can be made to generate conditionally by adding conditioning information $\boldsymbol{s}$ into the neural network function $f(\boldsymbol{z}_{t},t,\boldsymbol{s})$, e.g. implemented using a cross-attention in a U-Net (Ronneberger et al., 2015). When $\boldsymbol{s}$ sufficiently describes $\boldsymbol{x}$, using a diffusion model does not necessarily yield improvements compared to directly predicting $\boldsymbol{x}$ given $\boldsymbol{s}$. However, if $\boldsymbol{s}$ contains only partial information about $\boldsymbol{x}$ (such as a language description or a high-level scene description), then a conditional diffusion model will be able to utilize that information in addition to correlating pixels. Compared to directly predicting $\boldsymbol{x}$ given $\boldsymbol{s}$ in this scenario, a diffusion model will be able to generate high frequency details even if precise knowledge (such as about location, or shape) is not entirely given by $\boldsymbol{s}$. In contrast, directly predicting $\boldsymbol{x}$ from $\boldsymbol{s}$ (for instance after training with mean-squared error) would result in blurry predictions due to incomplete information.
@@ 3.1 Decoder Architecture & Conditioning
The DORSal decoder uses a convolutional U-Net architecture as is conventional in the diffusion literature (Ho et al., 2020). To attain consistency between $L$ views generated in parallel, following Video Diffusion (Ho et al., 2022c), each frame has feature maps which are enriched with 2d convolutions to process information within each frame and axial (self-)attention to propagate information between frames. We refer to this as a Multiview U-Net in our setting as each frame corresponds to a separate view of a scene. This is different from the X-UNet used in 3DiM (Watson et al., 2023), which relies on cross-attention between conditioning views and generated views, and therefore scales less favourably when increasing the number of output views $L$. DORSal relies on Object Slots for context about the scene, which avoids the cost of attending directly to large sets of conditioning features that are often redundant.
@@
We apply this conditioning in the same way that text is treated in recent work on text-to-image models Saharia et al. (2022c), i.e. integrated into the U-Net in two ways: 1) we attention-pool Radford et al. (2021) conditioning embeddings into a single embedding for modulating U-Net feature maps via FiLM (Perez et al., 2018), and 2) we use cross-attention (Vaswani et al., 2017) to attend on conditioning embeddings (keys) from the feature map (queries).
Independent slot conditioning.
Ideally, slot representations that summarize the scene should be conditionally independent given an image $\boldsymbol{z}_{t}$, $p(\boldsymbol{s}_{1:K}|\boldsymbol{z}_{t})=\prod_{k=1,K}p(\boldsymbol{s}_{i}|\boldsymbol{z}_{t})$, i.e. to be able to manipulate the presence of objects independently for editing purposes. In reality, the slot representations may respect this assumption to varying degrees, with an OSRT model trained with instance-level supervision (Sup-OSRT) being more likely to achieve this. However, even if slots would exclusively bind to particular regions of the encoded input views that correspond to individual objects, slots may still share information as the input view encoder has a global receptive field. To mitigate this issue, we consider dropping slots from the conditioning set independently following a Bernoulli rate set as a hyper-parameter $\lambda_{sd}$. The model thus sees slot subsets at training time (such that edits are now effectively in-distribution). As we will see, this allows for qualitatively better edits.
@@ 3.2 Editing & Sampling
At inference time, we explore a simple form of scene editing: by removing individual slots, we can—if the slot succinctly describes an individual object in the scene—remove that object from the scene. We remove slots by masking out the value of the slot, including any attention weights derived from it. Sampling with this edited conditioning yields $K$ edited scene renderings, where $K$ is the number of object slots in each model. We can then derive the effect of each edit by comparing it to unedited samples generated by keeping all slots for conditioning. To measure success, and to compare between methods, we propose to segment pixels based on whether they were affected by removing a particular slot, and compare to ground-truth instance segments using standard segmentation metrics.
@@
View-consistent sampling
Repeatedly generating blocks of $L$ frames is fast, but there is no guarantee on the consistency between the different blocks. This is because sampling from a conditional generative model inherently adds bits of information to the conditioning signal to produce a one-to-many mapping. This is useful in cases of partial observability, such as occlusion, wherein there exist multiple possibilities for unobserved scene content. However, this means that achieving consistency across views involves synchronizing the manner in which bits of information are added, which is challenging as the number of output views grows beyond the amount used during training (eg. $L=$ 3, 5 or 10). One approach is to alter the diffusion process to include a selection of already generated previous frames to generate the next set. This strategy is auto-regressive, and thus can be slower (only one additional frame is generated each time), and lead to exploding errors for longer auto-regressive chains.
Instead, we leverage the iterative nature of the generative denoising process to create smooth transitions as well as global consistency between frames. Our technique is inspired by Hoogeboom et al. (2023), where high resolution images are generated with overlapping patches by dividing the typical denoising process with $T$ steps is divided into multiple stages. To apply this technique for 3D camera-path rendering of up to 190 views in Section 5.3, we propose to interleave 3 types of frame shuffling for subsequent stages: 1) no shuffle (identity), to allow the model to make blocks of the context length consistent; 2) shift the frames in time by about half of the context length, which puts frames together with new neighbours in their context, allowing the model to create smooth transitions; 3) shuffle all frames with a random permutation, to allow the model to resolve inconsistencies globally.
@@ 4 Related Work
Motivated by NeRF Mildenhall et al. (2020), significant advances have recently been achieved in neural rendering Tewari et al. (2022). From many observations, NeRF optimizes an MLP through volumetric rendering, thereby allowing high-quality NVS. While several works extend this method to generalizing from few observations per scene Yu et al. (2021); Chen & Xu (2021), they do not provide accessible latent representations. Several latent 3D representation methods exist Sitzmann et al. (2019); Eslami et al. (2018); Moreno et al. (2023), however they do not scale beyond simple synthetic datasets. The recently proposed Scene Representation Transformer (SRT (Sajjadi et al., 2022c)) and extensions (RUST (Sajjadi et al., 2022b)) use large set-latent scene representations to scale to complex real-world datasets with or without pose information. However, SRT often produces blurry images due the L2-loss and high uncertainty in unobserved regions. While approaches like (Rombach et al., 2021) consider generative models for NVS, attaining 3d consistency is challenging with auto-regressive models.
Diffusion.
@@
Modern score-based diffusion models (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020) have been very successful in multiple domains. They learn to approximate a small step of a denoising process, the reverse of the pre-defined diffusion process. This setup has proven to be very successful and easy to use compared to other generative approaches such as variational autoencoders (Kingma et al., 2023), normalizing flows (Rezende & Mohamed, 2015) and adversarial networks (Goodfellow et al., 2014). Examples where diffusion models have had success are generation of images (Ho et al., 2022b; Dhariwal & Nichol, 2021), audio (Kong et al., 2021), and video (Ho et al., 2022a). Moreover, the extent to which they can be steered to be consistent with conditioning signals Ho & Salimans (2021); Nichol et al. (2021) has allowed for much more controllable image generation. More recently, pose-conditional image-to-image diffusion models have been applied to 3D NVS (Watson et al., 2023; Liu et al., 2023; Gu et al., 2023), focusing mainly on 3D synthesis of individual objects as opposed to complex visual scenes. DORSal leverages video diffusion models (Ho et al., 2022c) and object-slot conditioning to synthesize novel views that are more consistent, especially in real-world settings, and support object-level edits.
Object representations.
Object-centric methods for representation learning (Greff et al., 2020) aim at structuring the representation of, e.g., an image, a video, or of a 3D scene, in terms of its object content. This provides a direct interface for scene editing as well as object-centric downstream tasks, such as detection, tracking or instance segmentation. Most successful methods adopt a slot-based approach (Greff et al., 2019; Locatello et al., 2020; Singh et al., 2022; Seitzer et al., 2023), i.e. they learn a set of object slots to describe the content of a scene. In the context of 3D scene representation learning, ObSuRF (Stelzner et al., 2021), uORF (Yu et al., 2022), and OSRT (Sajjadi et al., 2022a) are recent examples of combining slot-based models with methods for 3D representation learning. We base DORSal on object slots obtained from OSRT as this method scales to large and diverse datasets. As explored in Slot-TTA (Prabhudesai et al., 2023), OSRT can further benefit from instance mask supervision, if available. In concurrent work, slot-based methods have also been explored in combination with diffusion-based decoders: LSD (Jiang et al., 2023) and SlotDiffusion (Wu et al., 2023) combine Slot Attention with a diffusion decoder in latent space for image and (for the latter) video object segmentation. Neither approach, however, considers 3D scenes or NVS, but solely focus on auto-encoding objectives.

### v3

@@ Abstract
Recent progress in 3D scene understanding enables scalable learning of representations across large datasets of diverse scenes. As a consequence, generalization to unseen scenes and objects, rendering novel views from just a single or a handful of input images, and controllable scene generation that supports editing, is now possible. However, training jointly on a large number of scenes typically compromises rendering quality when compared to single-scene optimized models such as NeRFs. In this paper, we leverage recent progress in diffusion models to equip 3D scene representation learning models with the ability to render high-fidelity novel views, while retaining benefits such as object-level scene editing to a large degree. In particular, we propose DORSal, which adapts a video diffusion architecture for 3D scene generation conditioned on frozen object-centric slot-based representations of scenes. On both complex synthetic multi-object scenes and on the real-world large-scale Street View dataset, we show that DORSal enables scalable neural rendering of 3D scenes with object-level editing and improves upon existing approaches.
Work done while interning at Google, equal contribution.
@@ 1 Introduction
In this work, we combine techniques from both of these subfields to further neural 3D scene rendering. We leverage frozen object-centric scene representations to condition probabilistic diffusion decoders capable of synthesizing novel views while also handling uncertainty about the scene. In particular, we use Object Scene Representation Transformer (OSRT) (Sajjadi et al., 2022a) to compute a set of Object Slots for a visual scene from only few observations, and condition a video diffusion architecture (Ho et al., 2022c) with these slots to generate sets of 3D consistent novel views of the same scene. We show that conditioning on object-level representations allows for scaling more gracefully to complex scenes, large sets of target views, and enables basic object-level scene editing by removing slots or by transferring them between scenes.
In summary, our contributions are as follows:
  • We introduce Diffusion for Object-centric Representations of Scenes et al. (DORSal), an approach to controllable 3D novel-view synthesis combining (frozen) object-centric scene representations with diffusion decoders.
  • Compared to prior methods from the 3D scene understanding literature (Sajjadi et al., 2022a; c), DORSal renders novel views that are significantly more precise (e.g. 5x-10x improvement in FID) while staying true to the content of the scene. Compared to prior work on 3D Diffusion Models (Watson et al., 2023), DORSal scales to more complex scenes, performing significantly better on real-world Street View data.
@@ 2 Preliminaries
Core to our approach to scene generation is the use of (pre-trained) object representations as conditioning information, as opposed to, e.g., conditioning on language prompts Ramesh et al. (2021); Rombach et al. (2022); Ho et al. (2022c). Recent breakthroughs in neural rendering have inspired multiple works for learning such 3D-centric object representations, including uORF (Yu et al., 2022) and ObSuRF (Stelzner et al., 2021). However, these methods do not scale beyond simple datasets due to the high memory and compute requirements of volumetric rendering. More recently, the Object Scene Representation Transformer (OSRT) (Sajjadi et al., 2022a) has been proposed as a powerful method that scales to much more complex datasets with wider camera pose distributions such as MultiShapeNet (Sajjadi et al., 2022c). Building upon SRT (Sajjadi et al., 2022c), it uses light-field rendering to obtain speed-ups by a factor of $\mathcal{O}(100)$ at inference time. We use OSRT as a base model for obtaining object representations as conditioning information for DORSal.
Figure 1: Model overview. (a) OSRT is trained to predict novel views through an Encoder-Decoder architecture with an Object Slot latent representation of the scene. Since the model is trained with the L2 loss and the task contains significant amounts of ambiguity, the predictions are commonly blurry. (b) After training the OSRT model, and freezing it, we take the Object Slots and combine it with the target Poses to be used as conditioning. Our Multiview U-Net is trained in a diffusion process to denoise novel views while cross-attending into the conditioning features (see Figure 2 for details). This results in sharp renders at test time, which can still be decomposed into the objects in the scene to support edits.
@@
Once the input views are encoded into the Object Slots, arbitrary novel views can be rendered by passing the target ray origin and direction (the Pose) into the Decoder. To encourage an object-centric decomposition in the Object Slots, Spatial Broadcast Decoders Watters et al. (2019) are commonly used in the literature: Each slot is decoded independently into a pair of RGB and alpha using the same decoder, after which a Softmax over the slots decides on the final output color. Since OSRT is trained end-to-end with the L2 loss, any uncertainty about novel views necessarily leads to blur in the final renders. OSRT can be trained fully unsupervised (in the absence of object labels) or using segmentation supervision (Prabhudesai et al., 2023) to guide the decomposition process.
@@
Denoising Diffusion Probabilistic Models (DDPMs) learn to generate data ${\bm{x}}$ by learning the reverse of a simple destruction process (Sohl-Dickstein et al., 2015). Such a diffusion process is convenient to express in its marginal form:
$q({\bm{z}}_{t}|{\bm{x}})=\mathcal{N}({\bm{z}}_{t}|\alpha_{t}{\bm{x}},\sigma_{t}^{2}\mathbf{I}),$ (1)
where $\alpha_{t}$ is a decreasing function and $\sigma_{t}$ is an increasing function over diffusion time $t\in[0,1]$. A neural network is then used to approximate $\bm{\epsilon}_{t}$, the reparametrization noise, to sample ${\bm{z}}_{t}$:
$L=\mathbb{E}_{t\sim\mathcal{U}(0,1),\bm{\epsilon}_{t}\sim\mathcal{N}(0,\mathbf{I})}\Big{[}w(t)||\bm{\epsilon}_{t}-f({\bm{z}}_{t},t)||^{2}\Big{]},$ (2)
where $f$ is a neural network and ${\bm{z}}_{t}=\alpha_{t}{\bm{x}}+\sigma_{t}\bm{\epsilon}_{t}$. There exists a particular weighting $w(t)$ for this objective to be a variational negative lower bound on $\log p({\bm{x}})$, although in practice the constant weighting $w(t)=1$ has been found to be superior for sample quality Ho et al. (2020); Kingma et al. (2021). Because diffusion models learn to correlate the pixels in their generations, they are able to generate images with crisp details even if the exact location of such details is not entirely known. We follow the framework of conditional diffusion models, where conditioning information ${\bm{s}}$, such as text or, in our case, information about scene content, is provided to the neural network function $f({\bm{z}}_{t},t,{\bm{s}})$, e.g. implemented using a cross-attention in a U-Net (Ronneberger et al., 2015).
@@ 3.1 Decoder Architecture & Conditioning
The DORSal decoder uses a convolutional U-Net architecture as is conventional in the diffusion literature (Ho et al., 2020). To attain consistency between $L$ views generated in parallel, following Video Diffusion (Ho et al., 2022c), each frame has feature maps which are enriched with 2d convolutions to process information within each frame and axial (self-)attention to propagate information between frames (see also Appendix B.2). We refer to this as a Multiview U-Net in our setting as each frame corresponds to a separate view of a scene. DORSal relies on Object Slots for context about the scene, which avoids the cost of attending directly to large sets of conditioning features that are often redundant.
@@
We apply this conditioning in the same way that text is treated in recent work on text-to-image models (Saharia et al., 2022c), i.e. integrated into the U-Net in two ways: 1) we attention-pool Radford et al. (2021) conditioning embeddings into a single embedding for modulating U-Net feature maps via FiLM (Perez et al., 2018), and 2) we use cross-attention (Vaswani et al., 2017) to attend on conditioning embeddings (keys) from the feature map (queries).
@@ 3.2 Editing & Sampling
At inference time, we explore a simple form of scene editing: by removing individual slots, we can—if the slot succinctly describes an individual object in the scene—remove that object from the scene. We remove slots by masking out the value of the slot, including any attention weights derived from it. Sampling with this edited conditioning yields $K$ edited scene renderings, where $K$ is the number of object slots in each model. We can then derive the effect of each edit by comparing it to unedited samples generated by keeping all slots for conditioning. To measure success, and to compare between methods, we propose to segment pixels based on whether they were affected by removing a particular slot, and compare to ground-truth instance segments using standard segmentation metrics. We further demonstrate successful transfer of objects between scenes as another form of scene editing.
@@
View-consistent camera path sampling.
Repeatedly generating blocks of $L$ frames is fast, but there is no guarantee on the consistency between the different blocks. This is because sampling from a conditional generative model inherently adds bits of information to the conditioning signal to produce a one-to-many mapping. Hence, achieving consistency across views involves synchronizing the manner in which bits of information are added, which is challenging as the number of output views grows beyond the amount used during training (as is required for generating long videos or smooth camera paths). We leverage the iterative nature of the generative denoising process to create smooth transitions as well as global consistency between frames. Our technique is inspired by Hoogeboom et al. (2023), where high resolution images are generated with overlapping patches by dividing the typical denoising process into multiple stages. For 3D camera-path rendering of hundreds of frames, we propose to interleave 3 types of frame shuffling for subsequent stages, while denoising only for a small number of steps per stage: 1) no shuffle (identity), to allow the model to make blocks of the context length consistent; 2) shift the frames in time by about half of the context length, which puts frames together with new neighbours in their context, allowing the model to create smooth transitions; 3) shuffle all frames with a random permutation, to allow the model to resolve inconsistencies globally.
@@ 4 Related Work
Motivated by NeRF Mildenhall et al. (2020), significant advances have recently been achieved in neural rendering Tewari et al. (2022). From many observations, NeRF optimizes an MLP through volumetric rendering, thereby allowing high-quality NVS. While several works extend this method to generalizing from few observations per scene Yu et al. (2021); Chen & Xu (2021), they do not provide accessible latent representations. Several latent 3D representation methods exist Sitzmann et al. (2019); Eslami et al. (2018); Moreno et al. (2023), however they do not scale beyond simple synthetic datasets. The recently proposed Scene Representation Transformer (SRT, Sajjadi et al. (2022c)) and extensions (RUST, Sajjadi et al. (2022b)) use large set-latent scene representations to scale to complex real-world datasets with or without pose information. However, SRT often produces blurry images due the L2-loss and high uncertainty in unobserved regions. While approaches like Rombach et al. (2021) consider generative models for NVS, attaining 3d consistency is challenging with auto-regressive models.
Diffusion Generative Models.
@@
Modern score-based diffusion models (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020) have been very successful in multiple domains. They learn to approximate a small step of a denoising process, the reverse of the pre-defined diffusion process. This setup has proven to be very successful and easy to use compared to other generative approaches such as variational autoencoders (Kingma et al., 2021), normalizing flows (Rezende & Mohamed, 2015) and adversarial networks (Goodfellow et al., 2014). Examples where diffusion models have had success are generation of images (Ho et al., 2022b; Dhariwal & Nichol, 2021), audio (Kong et al., 2021), and video (Ho et al., 2022a). Moreover, the extent to which they can be steered to be consistent with conditioning signals Ho & Salimans (2021); Nichol et al. (2022) has allowed for much more controllable image generation. More recently, pose-conditional image-to-image diffusion models have been applied to 3D NVS (Watson et al., 2023; Liu et al., 2023; Gu et al., 2023; Chan et al., 2023; Tewari et al., 2023), focusing mainly on 3D synthesis of individual objects as opposed to complex visual scenes. Chan et al. (2023) presents results for some indoor scenes, though it remains unclear how to manipulate the generated scenes at the object level. DORSal leverages video diffusion models (Ho et al., 2022c) and object-slot conditioning to synthesize novel views that are more consistent, especially in real-world settings, and support object-level edits.
Object-centric methods have also been explored in combination with diffusion-based decoders: LSD (Jiang et al., 2023) and SlotDiffusion (Wu et al., 2023) combine Slot Attention with a diffusion decoder in latent space for image and (for the latter) video object segmentation. Neither approach, however, considers 3D scenes or NVS, but solely focus on auto-encoding objectives. In concurrent work, OBJECT 3DIT (Michel et al., 2023) finetunes a 3D diffusion model (Liu et al., 2023) on supervised object edits obtained from synthetically generated data of scene edits. In contrast to this work, scene edits afforded by DORSal do not require supervision or specifically prepared data of scene edits. Alternatively, DisCoScene Xu et al. (2023), conditions adversarially-trained generators on an object-based scene layout to obtain a spatially disentangled generative radiance field from which a novel view can be rendered. In contrast, here we consider learned object representations as a conditioning signal, which further reduces the amount of prior knowledge needed about a scene.